OpenStudy (anonymous):

Five years ago a girl was one-half as old as her brother was then, and now she is three years younger than he is. What are their present ages? girl is 7; brother is 10 girl is 8; brother is 11 girl is 9; brother is 12

OpenStudy (anonymous):

a b or c

OpenStudy (anonymous):

Okay if you say that the girl is the variable g and the brother is the variable b, g-5=1/2(b-5) and g-3=b These two equations are given. System of equations time!

OpenStudy (anonymous):

ok